In an experiment, 57.4 g of iron (II) chloride solution reacts with 45.3 g of sodium hydroxide solution.

FeCla(an) + 3NaOH(an) -› Fe(OH) scs) + 3NaC|(ag)


Which reactant is the limiting reagent?


How much (in grams) of each product forms?


How much of the excess reagent (in grams) will be leftover after the reaction is complete?


1
Expert's answer
2022-05-13T03:48:03-0400

FeCl2 + 2NaOH = Fe(OH)2 + 2NaCl

n(FeCl2) = m(FeCl2) / Mr(FeCl2) = 57.4 g / 126.75 g/mol = 0.45 moles

n(NaOH) = m(NaOH) / 2Mr(NaOH) = 45.3 g / 2*40 g/mol = 0.57 moles


NaOH - excess reagent

FeCl2 - limiting reagent


m(Fe(OH)2) = (m(FeCl2) * Mr(Fe(OH)2) / Mr(FeCl2) = (57.4 g * 89.86 g/mol) / 126.75 g/mol = 40.69 g

m(NaCl) = (m(FeCl2) * 2Mr(NaCl) / Mr(FeCl2) = ( 57.4 g * (2*58.44 g/mol)) / 126.75 g/mol = 52.93 g


m(NaOH) = (m(FeCl2) * 2Mr(NaOH) / Mr(FeCl2) = ( 57.4 g * (2*40 g/mol)) / 126.75 g/mol = 36.23 g

m(NaOH) after the reaction is complete = 45.3 g - 36.23 g = 9.07 g